OpenStudy (anonymous):

what's the slope of (7,1)(8-5)

OpenStudy (karatechopper):

-6

OpenStudy (anonymous):

(y2 - y1) / (x2 - x1) = (-5 - 1) / (8 - 7) = (-6) / (1) = -6